OpenStudy (dj3strella):

Help me please i don't understand What is the GCF of 12x^3y^5 + 18x^4y^2

OpenStudy (solomonzelman):

\(\Large\color{green}{ \bf 12x^3y^5= 3 \times 2 \times 2 \times x^3 \times y^2 \times y^3}\) \(\Large\color{green}{ \bf 18x^4y^2 =3 \times 3 \times 2 \times x^3 \times x \times y^2}\) No take what they have in common and combine. \(\Large\color{green}{ \bf 12x^3y^5= \color{red} { 3 } \times \color{red} { 2 } \times 2 \times \color{red} { x^3 } \times \color{red} { y^2 } \times y^3}\) \(\Large\color{green}{ \bf 18x^4y^2 =3 \times \color{red} { 3 } \times \color{red} { 2 } \times \color{red} { x^3 } \times x \times \color{red} { y^2 } }\)

OpenStudy (solomonzelman):

I meant t say Now (nstead of "no") take what they have in common

OpenStudy (dj3strella):

ok thank you

OpenStudy (solomonzelman):

You are welcome... tell me what you get, just in case...

OpenStudy (dj3strella):

6x4y5 @SolomonZelman

OpenStudy (solomonzelman):

you don't have x^4 and y^5 in common, you only have x^3 and y^2 in common.

OpenStudy (dj3strella):

6x3y2

OpenStudy (solomonzelman):

yeah. I want to ask you, to avoid confusions, to use ^ \(\Large\color{blue}{ \sf x^{2} }\) would be written as x^2 and \(\Large\color{blue}{ \sf x~^{(v+2)} }\) would be written as x^(v+2) Okay?

OpenStudy (solomonzelman):

^ is just to say "raised to the power"